Freight Terms

Determine the amount to be paid in full settlement of each of two invoices (a) and (b), assuming that credit for returns and allowances was received prior to payment and that all invoices were paid within the discount period. If required, round the answers to the nearest dollar.

Merchandise Freight
Paid by Seller
Freight Terms Returns and
Allowances
a. $9,500 $300 FOB shipping point, 2/10, n/30 $1,600
b. 3,400 400 FOB destination, 1/10, n/30 1,000
a. $
b. $
Answer #1

Calculate amount to be paid

a) Amount to be paid = (Amount-return and allowance-Discount)+Freight = (9500-1600)*98%+300 = 8042

b) Amount to be paid = (Amount-return and allowance-Discount) = (3400-1000)*99% = 2376